# Global-Local Progressive Integration Network for Blind Image Quality Assessment

### v1

@@ V-A1 Databases and protocols
A comprehensive evaluation of the proposed model was conducted across ten IQA benchmark databases for IQA, including five synthetic distortion datasets, specifically LIVE [22], CSIQ [43], TID2013 [23], KADID-10k [24], and LIVE-MD [44], alongside five authentic distortion datasets, namely BID [45], CLIVE [46], KonIQ-10k [47], SPAQ [48], and LIVEFB [49]. Within the synthetic distortion datasets, the LIVE-MD dataset features images with multiple distortion types, whereas the remaining databases primarily focus on images with a single distortion type. Synthetic distortion datasets are constructed by algorithmically degrading original images, resulting in minimized content variation but a wide range of artificial distortions. In contrast, authentic distortion datasets comprise images captured in real-world settings, exhibiting substantial content diversity from varied scenes, objects and conditions encountered. The details of sample size, image resolution, and label types for each database are summarized in Table I.
Figure 7: The pre-training results of the proposed dataset.
The evaluation of BIQA methods employs two widely-adopted correlation metrics: Spearman’s rank-order correlation coefficient (SROCC) and Pearson’s linear correlation coefficient (PLCC). SROCC quantifies the monotonic association between predicted scores and ground truth, while PLCC measures their linear relationship. Specifically, SROCC computes the squared disparity between the ranks of predicted quality scores and MOS values. On the other hand, PLCC quantifies linear relationship between predicted quality scores and MOS values by calculating the covariance divided by the product of their standard deviations.
TABLE I: Comprehensive Summary of IQA Databases for Benchmarking BIQA Models.
Database Ref. Image Dist. Type Dist. Image Image Label Database
Count Count Count Resolution Type Type
LIVE [22] 29 5 779 480$\times$720 $\sim$ 768$\times$512 DMOS Synthetic
CSIQ [43] 30 6 866 512$\times$512 DMOS Synthetic
TID2013 [23] 25 24 3,000 384$\times$512 MOS Synthetic
KADID-10k [24] 81 25 10,125 384$\times$512 MOS Synthetic
LIVE-MD [44] 15 2 150 720$\times$1280 DMOS Synthetic
BID [45] - - 585 960$\times$1280 $\sim$ 1704$\times$2272 MOS Authentic
LIVEC [46] - - 1,162 500$\times$500 $\sim$ 640$\times$960 MOS Authentic
KonIQ-10k [47] - - 10,073 384$\times$512/768$\times$1024 MOS Authentic
SPAQ [48] - - 11,125 1080$\times$1440$\sim$6556$\times$3744 MOS Authentic
LIVEFB [49] - - 39,810 160$\times$186 $\sim$ 660$\times$1200 MOS Authentic
V-A2 Implementation details
@@
TABLE II: Benchmarking performance against existing BIQA models on authentic distortion databases, where the best and second-best performances are indicated by black bolded and gray bolded entries, respectively.
@@
TABLE III: Comparison to state-of-the-art BIQA models on five synthetic IQA databases, where the best and second-best performances are indicated by black bolded and gray bolded entries, respectively. GlintIQA* was fine-tuned on the respective dataset using the pre-trained model on the proposed dataset.
@@
V-B Comparison with BIQA Methods on Individual Databases
@@
To evaluate the proposed method, we conducted experiments on five authentic and five synthetic distortions datasets. For synthetic distortion datasets, pristine images were randomly partitioned into two subsets at an 8:2 ratio, with respective distorted images forming training and test sets to avoid content overlap. The authentic distortion dataset was partitioned into training and test sets in an 8:2 ratio. This process was repeated ten times with different random partitions, and the median SROCC and PLCC values were reported. As detailed in Tables II and III, we benchmarked our method against nine existing deep learning-based BIQA approaches, on both authentic and synthetic datasets. The results of the deep learning-based models were evaluated using the reported results in the original papers or replicated using their publicly released code.
V-B1 Authentic database evaluation
@@
As shown in Table II, the proposed GlintIQA demonstrates consistently superior performance across five authentic distortion datasets, achieving an average SROCC of 0.840 and PLCC of 0.867. This represents improvements of 1.94% and 1.40% over TempQT, and 0.60% and 0.46% over DEIQT, respectively. Specifically, on the CLIVE, GlintIQA achieves best results, outperforming DEIQT by 1.83% in SROCC and 1.45% in PLCC. Similarly, for the KonIQ-10k, the proposed method achieves the highest performance, with SROCC and PLCC gains of 1.63% and 1.28% over the second-best method. In the BID and SPAQ datasets, both GlintIQA and DEIQT reach best performance, with our method exhibiting a superior PLCC. Additionally, GlintIQA secures competitive results on the LIVEFB. Overall, the proposed GlintIQA exhibits superior and more stable performance compared to existing BIQA methods. The improvement of GlintIQA can be attributed to two key design elements: firstly, the method integrates both global coarse- and local fine-grained image features, enhancing its sensitivity to various distortions. Secondly, the inclusion of CNNs, specifically CLFE and SEM, significantly augments the model’s inductive capacity to learn representations of different distortions. This enhanced inductive capability is particularly advantageous for achieving superior performance on smaller datasets such as CLIVE and BID.
V-B2 Synthetic database evaluation
Table III summarizes the performance of BIQA models on five synthetic distortion datasets. The final two columns present the average performance across four of these datasets, excluding KADID-10k. GlintIQA* was fine-tuned on each dataset using a model pre-trained on our proposed dataset. The proposed GlintIQA* achieves the highest average SROCC and PLCC. Specifically, on the LIVE, most BIQA methods perform well, with GlintIQA achieving SROCC and PLCC values close to those of the top-performing DEIQT. For KADID-10k, GlintIQA surpasses DEIQT by 1.91% in SROCC and 1.80% in PLCC. GlintIQA attains the highest SROCC on LIVEMD and competitive results on CSIQ. Its performance on TID2013 is slightly lower than that of some methods, as the limited number of original images and the diversity of distortion types introduce variability in the results. However, the pre-trained GlintIQA* shows significant improvements on CSIQ and TID2013, with SROCC and PLCC gains of 3.08% and 2.85% on CSIQ, and 7.00% and 5.21% on TID2013 compared to GlintIQA. On LIVE, GlintIQA already demonstrates high performance, leaving little room for pre-training improvement. LIVEMD includes compound distortions, while the pre-trained model is based on the proposed dataset with single distortions, resulting in minimal pre-training gains on this dataset.
@@ V-B2 Synthetic database evaluation
TABLE IV: Summary of the statistical performance of the proposed method and existing BIQA methods.
@@
V-B3 Statistical significant test
V-C Comparison with BIQA Methods on Cross-dataset Testing
To rigorously evaluate the generalization abilities of BIQA models, we trained them on a complete database and subsequently assessed their performance on distinct, unseen databases. To ensure the reliability of results, each experiment was repeated ten times with randomized initializations, and the median of the resulting SROCC scores is presented. The experiments of our model were trained for 300 epochs and tested at 5 epoch intervals.
V-C1 Cross synthetic database evaluation
Table V presents results for six cross-synthetic database tests. The proposed GlintIQA and GlintIQA* models demonstrate exceptional performance across various cross-synthetic database tests, where GlintIQA achieves an average SROCC of 0.781, while GlintIQA* achieves an even higher average SROCC of 0.839. In particular, GlintIQA surpasses Hu et al.’s method [32] by 3.64% and 13.70% when trained on LIVE and tested on CSIQ and TID2013, respectively. When trained on CSIQ, GlintIQA outperforms TReS by 0.64% on LIVE and the second-ranked method by 7.58% on TID2013. Trained on TID2013 and tested on LIVE, GlintIQA outperforms DB-CNN by 5.05%, and DB-CNN outperforms GlintIQA by 7.74% on CSIQ, attributed to additional data used for pre-training. However, GlintIQA* with extra pre-training data achieves an 11.30% improvement over DB-CNN.
Furthermore, GlintIQA* demonstrates advancements in several other cross-dataset tests, such as an 8.53% improvement over GlintIQA when trained on LIVE and tested on CSIQ, and a remarkable 16.60% improvement when trained on CSIQ and tested on TID2013. The results highlight the effectiveness of our proposed dataset for enhancing model generalization by providing diverse image content for the model to learn distortion and quality variations. In addition, GlintIQA stands out even without pre-training, surpassing existing methods thanks to its dual focus on global and local features for capturing image distortions. Its combination of convolutional networks and ViT further bolsters its ability to learn and generalize effectively.
@@ V-C1 Cross synthetic database evaluation
V-C2 Cross authentic database evaluation
Table VI presents the performance of six cross-authentic dataset tests. The proposed GlintIQA demonstrates the best results across all six cross-authentic database tests. GlintIQA achieves the highest overall average performance, surpassing the second-ranked DEIQT by a significant 5.04%. When trained on CLIVE, GlintIQA excels on both BID and KonIQ-10k, outperforming TReS by 2.79% and HyperIQA by 2.85%, respectively. GlintIQA, when trained on KonIQ-10k, surpasses DEIQT by 1.06% on BID and REQA by 3.12% on CLIVE. Notably, GlintIQA achieves even more impressive gains on the relatively small-scale BID dataset, outperforming the second-ranked methods in CLIVE and KonIQ-10k by 6.35% and 4.92%, respectively. These results provide convincing evidence of GlintIQA’s superior generalization capabilities, highlighting its ability to adapt effectively to diverse image quality assessment scenarios.
V-D Ablation Study
V-D1 Model components
To validate the effectiveness of individual components in the proposed GlintIQA, we conduct two types of ablation studies: evaluations on a single dataset and across-dataset tests. The model configurations are illustrated in Fig. 8 (omitting feature concatenation, reshaping, and flattening operations for clarity). The structures are described as follows: a) Utilize CLFE module to extract local multi-scale features and perform regression using MLP on the embedded features. b) Employ VGFE module to extract global multi-scale features and regress them with MLP. c) Jointly leverage both CLFE and VGFE modules to extract local and global multi-scale features followed by regression. d) The proposed model sans the SEM modules. e) The model in its complete form, as shown in Fig. 1.
@@ V-D1 Model components
Based on the performance comparisons in Tables VII and VIII, where the blue-highlighted values signify performance gains achieved by other model configurations relative to model a, several key observations can be made: 1) The model b employing VGFE, outperforms the model utilizing LFE by 3.42% in terms of SROCC on individual datasets, and 5.48% on cross-dataset evaluations. The results indicate that global features are capable of capturing more comprehensive information pertaining to image quality in comparison to local features. 2) Model c, which combines CLFE and VGFE, achieves consistent enhancement over model a, with a 3.75% gain in SROCC on individual datasets and a 5.78% gain on cross-dataset evaluations. These results demonstrate the complementary effects of local and global features. 3) Model d, which incorporates CWSA, attains a 3.96% improvement in SROCC on individual datasets and a 6.45% improvement on cross-dataset evaluations. This outcome suggests that CWSA can facilitate the model’s ability to better concentrate on the salient features of images. 4) Model e, which combines CWSA and SEM, achieves the best performance, with a 4.16% increment in SROCC on individual datasets and an 8.04% increment on cross-dataset evaluations. This suggests that the interactive stacking of CWSA and SEM is highly effective in improving the integration and utilization of multi-scale features, allowing an accurate evaluation of image quality.
@@
V-D2 The proposed dataset
To validate the effectiveness of our proposed dataset, besides the initial validations in Tables III and V, we also conducted cross-dataset tests on KADID-10k, pitting GlintIQA against two other BIQA models, i.e., TReS and DEIQT. Their experimental results were replicated utilizing the publicly accessible code under the same training and test sets. As shown in Table IX, the results shown that even without pretraining, GlintIQA confidently surpasses both TReS and DEIQT in these evaluations. Furthermore, the pretrained GlintIQA* amplifies these advantages, boasting average SROCC and PLCC improvements of 5.42% and 5.72% over its non-pre-trained counterpart, GlintIQA. Notably, our proposed dataset is constructed using the labels of KADID-10k. The performance gains achieved by the pre-trained model can be attributed to two key factors: our dataset sufficiently increases the diversity of image content and the effectiveness of the image quality label annotation method.
To facilitate visual comparison, we additionally conducted a group maximum differentiation (gMAD) competition [57] between GlintIQA and GlintIQA* on Waterloo database [58]. In the evaluation, a defender model categorizes image quality into six levels, while an attacker model aims to expose its weaknesses by identifying image pairs with significant quality discrepancies within those levels. Specifically, the attacker selects images that the defender model rates similarly but that the attacker model perceives as having a large quality gap. This approach reveals differences in quality perception between models. Fig. 9 shows the gMAD competition between GlintIQA and pretrained GlintIQA*. When fixed as the defender GlintIQA, GlintIQA* successfully recognizes perceptually different image pairs, whereas the attack by GlintIQA results in image pairs with negligible quality differences. This verifies that pretrained GlintIQA* can further enhance model generalizability, solidifying the validity of our proposed dataset.
TABLE IX: Ablation study of the proposed dataset on cross-database tests and comparison of BIQA models.

### v2

@@ IV-A1 Databases and protocols
A comprehensive evaluation of the proposed model was conducted across ten IQA benchmark databases, including five synthetic distortion datasets, specifically LIVE [28], CSIQ [33], TID2013 [29], KADID-10k [30], and LIVE-MD [34], alongside five authentic distortion datasets, namely BID [37], CLIVE [38], KonIQ-10k [39], SPAQ [40], and LIVEFB [41]. Within the synthetic distortion datasets, the LIVE-MD dataset features images with multiple distortion types, whereas the remaining databases primarily focus on images with a single distortion type. Synthetic distortion datasets are constructed by algorithmically degrading original images, resulting in minimized content variation but a wide range of artificial distortions. In contrast, authentic distortion datasets comprise images captured in real-world settings, exhibiting substantial content diversity from varied scenes, objects and conditions encountered. The details of sample size, image resolution, and label types for each database are summarized in Table I. The evaluation of NR-IQA methods employs two widely-adopted correlation metrics: Spearman’s rank-order correlation coefficient (SROCC) and Pearson’s linear correlation coefficient (PLCC). SROCC quantifies the monotonic association between predicted scores and ground truth, while PLCC measures their linear relationship.
@@
TABLE II: Benchmarking performance against existing NR-IQA models on authentic distortion databases, where the best and second-best performances are indicated by black bolded and gray bolded entries, respectively.
@@
TABLE III: Comparison to state-of-the-art NR-IQA models on five synthetic IQA databases, where the best and second-best performances are indicated by black bolded and gray bolded entries, respectively. GlintIQA* was fine-tuned on the respective dataset using the pre-trained model on the proposed SAQT-IQA dataset.
@@
IV-A2 Implementation details
IV-B Comparison with NR-IQA Methods on Individual Datasets
@@
To evaluate the proposed method, we conducted experiments on five authentic and five synthetic distortions datasets. For synthetic distortion datasets, pristine images were randomly partitioned into two subsets at an 8:2 ratio, with respective distorted images forming training and test sets to avoid content overlap. The authentic distortion dataset was partitioned into training and test sets in an 8:2 ratio. This process was repeated ten times with different random partitions, and the median SROCC and PLCC values were reported. As detailed in Tables II and III, we benchmarked our method against nine existing deep learning-based NR-IQA approaches, on both authentic and synthetic datasets. The results of the deep learning-based models were evaluated using the reported results in the original papers or replicated using their publicly released code.
IV-B1 Authentic database evaluation
@@
As shown in Table II, the proposed GlintIQA demonstrates consistently superior performance across five authentic distortion datasets, achieving an average SROCC of 0.840 and PLCC of 0.867. This represents improvements of 1.94% and 1.40% over TempQT, and 0.60% and 0.46% over DEIQT, respectively. Specifically, on the CLIVE, GlintIQA achieves best results, outperforming DEIQT by 1.83% in SROCC and 1.45% in PLCC. Similarly, for the KonIQ-10k, the proposed method achieves the highest performance, with SROCC and PLCC gains of 1.63% and 1.28% over the second-best method. In the BID and SPAQ datasets, both GlintIQA and DEIQT reach best performance, with our method exhibiting a superior PLCC. Additionally, GlintIQA secures competitive results on the LIVEFB. Overall, the proposed GlintIQA exhibits superior and more stable performance compared to existing NR-IQA methods. The improvement of GlintIQA can be attributed to two key design elements: firstly, the method integrates both global coarse- and local fine-grained image features, enhancing its sensitivity to various distortions. Secondly, the inclusion of CNNs, specifically CLFE and SIEM, significantly augments the model’s inductive capacity to learn representations of different distortions. This enhanced inductive capability is particularly advantageous for achieving superior performance on smaller datasets, e.g., CLIVE and BID.
@@
TABLE IV: Summary of the statistical performance of the proposed method and existing NR-IQA methods.
@@
IV-B2 Synthetic database evaluation
Table III summarizes the performance of NR-IQA models on five synthetic distortion datasets. The final two columns present the average performance across four of these datasets, excluding KADID-10k. GlintIQA* was fine-tuned on each dataset using a model pre-trained on our proposed dataset. The proposed GlintIQA* achieves the highest average SROCC and PLCC. Specifically, on the LIVE, most NR-IQA methods perform well, with GlintIQA achieving SROCC and PLCC values close to those of the top-performing DEIQT. For KADID-10k, GlintIQA surpasses DEIQT by 1.91% in SROCC and 1.80% in PLCC. GlintIQA attains the highest SROCC on LIVEMD and competitive results on CSIQ. Its performance on TID2013 is slightly lower than that of some methods, as the limited number of original images and the diversity of distortion types introduce variability in the results. However, the pre-trained GlintIQA* shows significant improvements on CSIQ and TID2013, with SROCC and PLCC gains of 3.08% and 2.85% on CSIQ, and 7.00% and 5.21% on TID2013 compared to GlintIQA. On LIVE, GlintIQA already demonstrates high performance, leaving little room for pre-training improvement. LIVEMD includes compound distortions, while the pre-trained model is based on the proposed dataset with single distortions, resulting in minimal pre-training gains on this dataset.
@@ IV-B2 Synthetic database evaluation
IV-B3 Statistical significant test
IV-C Comparison with NR-IQA Methods on Cross-dataset Tests
To rigorously evaluate the generalization abilities of NR-IQA models, we trained them on a complete database and subsequently assessed their performance on distinct, unseen databases. To ensure the reliability of results, each experiment was repeated ten times with randomized initializations, and the median of the resulting SROCC scores is presented. The experiments of our model were trained for 300 epochs and tested at 5 epoch intervals.
IV-C1 Cross synthetic database evaluation
Table V presents results for six cross-synthetic database tests. The proposed GlintIQA and GlintIQA* models demonstrate exceptional performance across various cross-synthetic database tests, where GlintIQA achieves an average SROCC of 0.781, while GlintIQA* achieves an even higher average SROCC of 0.839. In particular, GlintIQA surpasses Hu et al.’s method [49] by 3.64% and 13.70% when trained on LIVE and tested on CSIQ and TID2013, respectively. When trained on CSIQ, GlintIQA outperforms TReS by 0.64% on LIVE and the second-ranked method by 7.58% on TID2013. Trained on TID2013 and tested on LIVE, GlintIQA outperforms DB-CNN by 5.05%, and DB-CNN outperforms GlintIQA by 7.74% on CSIQ, attributed to additional data used for pre-training. However, GlintIQA* with extra pre-training data achieves an 11.30% improvement over DB-CNN. Furthermore, GlintIQA* demonstrates advancements in several other cross-dataset tests, e.g., an 8.53% improvement over GlintIQA when trained on LIVE and tested on CSIQ, and a remarkable 16.60% improvement when trained on CSIQ and tested on TID2013. The results highlight the effectiveness of our proposed dataset for enhancing model generalization by providing diverse image content for the model to learn distortion and quality variations.
Figure 8: Ablation Study Results on feature integation strategies.
@@ IV-C1 Cross synthetic database evaluation
IV-C2 Cross authentic database evaluation
Table VI presents the performance of six cross-authentic dataset tests. The proposed GlintIQA demonstrates the best results across all six cross-authentic database tests. GlintIQA achieves the highest overall average performance, surpassing the second-ranked DEIQT by a significant 5.04%. When trained on CLIVE, GlintIQA excels on both BID and KonIQ-10k, outperforming TReS by 2.79% and HyperIQA by 2.85%, respectively. GlintIQA, when trained on KonIQ-10k, surpasses DEIQT by 1.06% on BID and REQA by 3.12% on CLIVE. Notably, GlintIQA achieves even more impressive gains on the relatively small-scale BID dataset, outperforming the second-ranked methods in CLIVE and KonIQ-10k by 6.35% and 4.92%, respectively. These exceptional results across both synthetic and authentic database evaluations demonstrate GlintIQA’s superior generalization capabilities, attributable to its dual focus on global and local features for capturing image distortions. The combination of convolutional networks and ViT further enhances its ability to learn and generalize effectively across diverse image quality assessment scenarios.
@@ IV-C2 Cross authentic database evaluation
IV-D Ablation Study
IV-D1 Integation strategy
In this ablation study, we explore two feature fusion strategies for integrating multi-scale features from CLFE and VGFE. “CLFE-to-VGFE”: This strategy starts with the initial fusion of CLFE features, followed by a progressive integration with VGFE features. As illustrated in Fig. 4, it emphasizes the importance of global representations while using local details as a supportive element. “VGFE-to-CLFE”: Conversely, this strategy begins with the fusion of VGFE features, followed by a progressive integration of CLFE features. As shown in Fig. 8, the “CLFE-to-VGFE” strategy consistently outperforms the “VGFE-to-CLFE” strategy across most datasets. The results underscore the superior importance of global features in quality assessment. Furthermore, it demonstrates that local details effectively complement VGFE’s global representations, enhancing overall performance. Nevertheless, both strategies achieve relatively high performance, demonstrating their viability in feature integration.
IV-D2 Model components
To validate the effectiveness of individual components in the proposed GlintIQA, we conduct two types of ablation studies: evaluations on a single dataset and across-dataset tests. The model configurations are illustrated in Fig. 9 (omitting feature concatenation, reshaping, and flattening operations for clarity). The structures are described as follows: a) Utilize CLFE module to extract local multi-scale features and perform regression using MLP on the embedded features. b) Employ VGFE module to extract global multi-scale features and regress them with MLP. c) Jointly leverage both CLFE and VGFE modules to extract local and global multi-scale features followed by regression. d) The proposed model sans the SIEM modules. e) The model in its complete form, as shown in Fig. 4.
Based on the performance comparisons in Tables VII and VIII, where the brown-highlighted values indicate performance improvements over model a), we make the following observations: Model b) with VGFE surpasses the CLFE-based model by 3.42% in SROCC on individual datasets and 5.48% on cross-dataset evaluations, demonstrating the superiority of global features in capturing comprehensive quality information. Model c), combining CLFE and VGFE, shows consistent improvements with 3.75% and 5.78% gains in SROCC for individual and cross-dataset evaluations respectively, verifying the complementary nature of local and global features. Model d) with CWSA achieves 3.96% and 6.45% improvements in SROCC for respective evaluations, indicating enhanced feature saliency detection. Model e), integrating both CWSA and SIEM, exhibits optimal performance with 4.16% and 8.04% SROCC improvements respectively, demonstrating the effectiveness of their interactive combination in multi-scale feature integration for accurate evaluation of quality assessment.
@@ IV-D2 Model components
IV-D3 The proposed SAQT-IQA dataset
To validate the effectiveness of our proposed dataset, besides the initial validations in Tables III and V, we also conducted cross-dataset tests on KADID-10k, pitting GlintIQA against two other NR-IQA models, i.e., TReS and DEIQT. Their experimental results were replicated utilizing the publicly accessible code under the same training and test sets. As shown in Table IX, the results shown that even without pretraining, GlintIQA confidently surpasses both TReS and DEIQT in these evaluations. Furthermore, the pretrained GlintIQA* amplifies these advantages, boasting average SROCC and PLCC improvements of 5.42% and 5.72% over its non-pre-trained counterpart, GlintIQA. Notably, our proposed dataset is constructed using the labels of KADID-10k. The performance gains achieved by the pre-trained model can be attributed to two key factors: our dataset sufficiently increases the diversity of image content and the effectiveness of the image quality label annotation method.
To facilitate visual comparison, we further conducted a group maximum differentiation (gMAD) competition [51] between GlintIQA and GlintIQA* on Waterloo database [52]. In the evaluation, a defender model categorizes image quality into six levels, while an attacker model aims to expose its weaknesses by identifying image pairs with significant quality discrepancies within those levels. Specifically, the attacker selects images that the defender model rates similarly but that the attacker model perceives as having a large quality gap. This approach reveals differences in quality perception between models. Fig. 10 shows the gMAD competition between GlintIQA and pretrained GlintIQA*. When fixed as the defender GlintIQA, GlintIQA* successfully recognizes perceptually different image pairs, whereas the attack by GlintIQA results in image pairs with negligible quality differences. This verifies that pretrained GlintIQA* can enhance model generalizability, solidifying the validity of our proposed dataset.
TABLE IX: Ablation study of the proposed SAQT-IQA dataset on cross-database tests and comparison of NR-IQA models.